Help me translate this into a quadratic equation and solve. Thanks!
A rectangle parking lot is 50 ft longer than its width. Determine the dimensions of the parking lot if it is measures 250 ft. diagonally.

Answer by scott8148(6628)

let x=width, so x+50=length

by Pythagoras __ x^2+(x+50)^2=250^2 __ x^2+x^2+100x+2500=62500 __ dividing by 2 __ x^2+50x+1250=31250

subtracting 31250 __ x^2+50x-30000=0 __ factoring __ (x+200)(x-150)=0

x+200=0 __ x=-200 __ negative value not realistic

x-150=0 __ x=150

150-200-250 is a 3-4-5 triangle